Is It Possible to have Two rectangles with the same area, but a side of one is twice as long as a side of the other? if yes then please draw an explain.
Found 2 solutions by ankor@dixie-net.com, greenestamps:
Answer by ankor@dixie-net.com(22740)

Is It Possible to have Two rectangles with the same area, but a side of one is twice as long as a side of the other? if yes then please draw an explain.
:
The area is the product of length and width,
If you double the length, you have to halve the width
for example: 8 * 6 = 48, 16 * 3 = 48, 32 * 1.5 = 48

Answer by greenestamps(13215)



A method that can often be used to simplify the process of multiplying two numbers is to double one of them and cut the other in half, keeping the product the same.

For example, if I had to multiply 35*18 mentally, I would double the 35 and cut the 18 in half, giving me 70*9. That's a lot easier than 35 times 18, because of the final "0" on the first number. 7 times 9 is 63; add the 0 to get the answer of 630.

So, to answer your question, you can take ANY rectangle, cut it in half and rearrange the two pieces to get a new rectangle with the same area that has its length twice the original length.

I can't put a drawing in my response; but I can describe the process.

Using my numerical example above, suppose you have a rectangle that is 35 long and 18 wide. Cut it in half lengthwise to get two rectangles each 35 long and 9 wide; now place the two halves end-to-end to get a new rectangle that is 70 long and 9 wide.

The new rectangle has the same area as the original, and one of its sides is twice the length of one of the sides of the original.